6Na + Fe2O3 --> 3Na2O + 2Fe If you are provided 100g Na and 100g of Fe3O2


1
Expert's answer
2021-01-25T07:01:44-0500

6Na +  Fe2O3 --> 3Na2O + 2Fe


"n(Na)=\frac{m}{M}=\frac{100}{22.9898}=4.35"


"n(Fe2O3)=\frac{m}{M}=\frac{100}{159.691}=0.63"


 Fe2O3 - limiting reagent

"m (Na2O)=n\times M=0.63\times61.979=38.81"

"m (Fe)=n\times M=0.63\times55.847=35.18"